OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#6846 — Marseille : BEA13 BRA13 COR13 DRA13 FER13 GAR13 MEN13 PRA13 SMG13 SML13 STM13
Incident Report for Web Cloud
Resolved
Following to a fiber interruption, we have just lost the following NRAs:

BEA13
BRA13
COR13
DRA13
FER13
GAR13
MEN13
PRA13
SMG13
SML13
STM13

A technician is working to audit the optical link.


Update(s):

Date: 2012-06-06 20:20:32 UTC
The fibers are welded.
The NRA are operational.

Following to the new weld we move from -18.4 to -19db

All websites are UP.


Date: 2012-06-06 20:16:04 UTC
Welds are being processed.

Date: 2012-06-06 20:15:06 UTC
SFR has finally found an interruption on the fiber.
Technicians are on site since 5pm.

The cable was partially defective.
It was cut a 5pm50 by the teams on site.

The cable is currently being prepared for welding.

The beginning of welds is expedted between 7:30pm and 8pm.

It is a 144 FO to weld.


Date: 2012-06-06 15:02:32 UTC
The optical link is operational in the sense JOL13 -> FER13. whereas, the light is turned off in the direction FER13 -> JOL13. We expect an FT audit in the dispatchers rooms.

Date: 2012-06-06 13:47:49 UTC
The interruption is not yet localized. We seek the origin of the interruption on the fiber.


Date: 2012-06-06 13:22:07 UTC
The technician is on the first site.
We check the links.

Date: 2012-06-06 13:21:16 UTC
The technician will be on the site at 14h.
Posted Jun 06, 2012 - 13:20 UTC